the areas of similar triangles are 25 and 16. The perimeter oh the first is 15, find the second perimeter.
I am confused on how to solve this. Every time I get some weird number that isn't the answer.

Answer by sachi(548)

In two similar triangles, the ratio of their areas is the square of the ratio of their sides =25/16
in Similar Triangles - ratios of parts, the perimeter, sides, altitudes and medians are all in the same ratio. Therefore, the area ratio will be the square of any of these ratios too.
so the ratio of the perimeters=sqrt(25/16)=5/4
one perimeter=15
so the 2nd perimeter=15*4/5=12
ans